OpenStudy (anonymous):

Solve 4 over x minus 4 equals the quantity of x over x minus 4, minus four thirds for x and determine if the solution is extraneous or not. x = 4, extraneous x = 4, non-extraneous x = -4, extraneous x = -4, non-extraneous

OpenStudy (anonymous):

\[\frac{4}{x-4}=\frac{x}{x-4}-\frac{4}{3}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

there is no solution, so not sure why you are supposed to come up with one

OpenStudy (anonymous):

i guess since there is no solution you are supposed to say x = 4, extraneous

OpenStudy (anonymous):

if you subtract you get \[\frac{4-x}{x-4}=-\frac{4}{3}\]

OpenStudy (anonymous):

since \(\frac{4-x}{x-4}=-1\) you have \[-1=-\frac{4}{3}\] which is silly

OpenStudy (anonymous):

not sure why you are supposed to say "4, extraneous" when "no solution" should suffice, but what do i know? must be some math teacher way to solve that gives \(4\) as an answer, even though it isn't one

OpenStudy (anonymous):

im not understanding

OpenStudy (anonymous):

subtract \(\frac{x}{x-4}\) from both sides of the equation

OpenStudy (phi):

how do people come up with these questions??!

OpenStudy (anonymous):

you get \[\frac{4-x}{x-4}=-\frac{3}{4}\]

OpenStudy (anonymous):

the left hand side is a number, it is \(-1\)

OpenStudy (anonymous):

the answer choices are under the question

OpenStudy (anonymous):

@phi no clue

OpenStudy (anonymous):

go with x = 4, extraneous

OpenStudy (anonymous):

extraneous is math teacherese for "no solution"

OpenStudy (anonymous):

ok ill let you know if its right or wrong i have 51 questions so this is just the many ones XD

OpenStudy (anonymous):

51 in 60 minutes lol

OpenStudy (anonymous):

huh

OpenStudy (phi):

if you clear the denominator (multiply by x-4 ) you can "solve" to get x=4 but really sats way of solving shows no solution with no extraneous value.

OpenStudy (mathstudent55):

\(\large \dfrac{4}{x-4}=\dfrac{x}{x-4}-\dfrac{4}{3} \) \(\large 3(x - 4)\dfrac{4}{x-4}=3(x - 4)\dfrac{x}{x-4}-3(x - 4)\dfrac{4}{3} \) \(\large 12 = 3x -12x + 48\) \(\large 9x = 36\) \(\large x = 4\) Since x = 4 cause a zero in the denominator, x = 4 is discarded and there is no solution.

OpenStudy (anonymous):

ok thank you <3